NOTE These are all part of Question 4

Feather Friends, Inc., distributes a high-quality wooden birdhouse that sells for $80 per unit. Variable expenses are $40.00Complete this question by entering your answers in the tabs below. Req 1 Req 2 Req 3 Req 4A Req 4B Req 5A Req 5B Req 6 HUUUUUComplete this question by entering your answers in the tabs below. Req 1 Req 2 Req 3 Req 4A Req 4B Req 5A Req 5B Req 6 Use thComplete this question by entering your answers in the tabs below. Req 1 Req 2 Req 3 Req 4A Req 4B Req 5A Req 5B Req 6 If thiComplete this question by entering your answers in the tabs below. Req 1 Req 2 Req 3 Req 4A Req 4B Req 5A Req 5B Req 6 ------Complete this question by entering your answers in the tabs below. Req 1 Req 2 Req 3 Req 4A Req 4B E Req 5A Req 5B Req 6 AssuComplete this question by entering your answers in the tabs below. Req 1 Req 2 Req 3 Req 4A Req 4B Req 5A Req 5B Req 6 ------Complete this question by entering your answers in the tabs below. Req 1 Req 2 Req 3 Req 4A Req 4B Req 5A Req 5B Req 6 The saComplete this question by entering your answers in the tabs below. Req 1 Req 2 Req3 Req 4A Req 4B Req 5A Req 5B Req 6 -------
Answer #1

Contribution Margin Ratio = Contribution Margin/Sales

= 1,040,000/2,080,000

= 50%

Break even point in sales dollars = Fixed costs/Contribution Margin Ratio

= 160,000/50%

= $320,000

Increase in operating income = increase in contribution margin since fixed costs remain same

= 50,000*50%

= $25,000

4-a Degree of Operating leverage = Contribution Margin/Net Operating income

= 1,040,000/880,000

= 1.1818

4-b % increase in Income = % increase in sales*Degree of operating leverage

= 12%*1.1818 = 14.1816%

i.e. 14.18%

5a. Net Operating income = (70.4 – 40)*26000*1.25 – 160,000-60,000

= $768,000

b.Decrease in Net income = -$112,000

6.Desired Operating Income = $880,000

Add: Fixed costs other than advertising= 160,000

Total = $1,040,000

Contribution Margin = (80-40-2.5)*26,000*1.25 = $1,218,750

Increase in advertising expense = 1,218,750-1,040,000 = $178,750